# Multilingual hierarchical classification of job advertisements for job vacancy statistics

### v1

@@ 4.3.1 The 10,000 ads dataset
Before sampling, the data was cleaned to remove rows that had no job description (233,464 observations), had zero words or less than 5 characters1818endnote: 18The number of words was counted using stri_count_words function from the stringi package and number of characters were counted using nchar from the base R (93 observations). Then we removed records with non-Polish descriptions. We have used Google’s Compact Language Detector 3 available through the cld3 package in R (Ooms,, 2024) to detect the language of the job description and identified 181,425 rows with non-Polish text. In the end, the final dataset used for sampling consisted of 1,390,985 records (77% of the initial dataset).
@@
The second stratification variable was based on the number of words in the description that were categorized into the following groups: $[1,10]$, $(10,25]$, $(25,50]$, $(50,100]$, $(100,150]$, $(150,200]$, $(200,300]$ and $300+$. Table 14 presents information about stratification and population sizes within these strata. The largest strata was "specjaliści" (specialists) with over 130k ads. The second largest were "sprzedawcy" (sellers) with close to 55k ads. The smallest group were "sprzątacze" (cleaners) with over 7k ads. According to the second stratification variable groups with over 50 words were more or less of the same size. There were about 40k ads with at most 10 words and 70k with at most 25 words.
In order to code the dataset, three experts from PEOs were recruited in February 2022. Due to budgetary limitations, the sample size was set at 10,000, with each expert responsible for coding 3,333 advertisements. However, due to rounding, the final sample size was 10,002. The sample size was selected using a stratified systematic sampling design, with probabilities proportional to the number of characters in a given ad. The sample was selected using strata from the sampling package (Tillé and Matei,, 2023). Further details regarding the sample size according to the source can be found in Table 15. The primary source was pracuj.pl, praca.egospodarka and aplikuj.pl. It should be noted that 241 ads were from CBOP, which enabled an assessment of the quality of CBOP encoding.
A total of 3,233 advertisements were selected for each expert, with an additional 100 advertisements sampled for cross-validation of expert coding1919endnote: 19it should be noted that for 4 ads we also had information from CBOP. Each expert was given an Excel spreadsheet, available via Google Sheets, containing all the relevant information. The experts were tasked with coding the advertisements into six-digit codes and providing comments where necessary. They were given a one-month period to complete the coding of the advertisements into occupation codes.
@@ 4.3.2 The 1,000 ads dataset
In January 2023, an additional dataset of 1,000 advertisements was selected for manual coding by a single expert. The sample was selected from a comprehensive dataset of over 543,000 ads scraped from one of the largest services in Poland. In the case of this particular dataset, we were able to identify in excess of 227,000 unique descriptions. We then proceeded to implement a sample selection process in accordance with the following scheme: the specification of a search phrase (including regular expression notation) and the subsequent selection of the most appropriate samples. The information pertaining to the aforementioned phrases, the selected advertisements and the number of cases in one of the largest services in Poland is detailed in Table 16.
@@ 4.4 The multilingual dataset
Firstly, the Polish text was translated into English using Google Translate. This was achieved by utilising the Google Sheets application and the googletranslate() function. This approach was employed for the translation of over 200,000 examples. Additionally, Argos Translate was employed for the translation of the remaining text. The open-source offline translation library, written in Python, is available at https://github.com/argosopentech/argos-translate. It uses the OpenNMT library (Klein et al.,, 2017), which in turn uses the CTranslate2 library, which is a fast inference engine for Transformer models. The objective was to evaluate the influence of an automated language library on the classification of a bilingual dataset.
@@ 5.1.1 The 1,000 dataset
The 1,000 dataset was coded by a single expert, and thus, cross-validation was not available. The expert identified 968 advertisements for which a single occupation code was deemed appropriate, while 32 advertisements were assigned multiple codes (27 with two codes, three with three codes, and one with four codes). Therefore, the final dataset comprised 1,035 advertisements with 226 codes, with the most prevalent being: "Programista aplikacji" (Applications programmer, 125 occurrences; 251401), "Analityk baz danych (data scientist)" (Data scientist, 77 occurrences; 252102) and "Specjalista do spraw cyberbezpieczeństwa" (Cybersecurity specialist, 63 occurrences; 252902). See Table 17 for the information on the number of unique 6-digit and 1-digit codes for a given stratum defined by the search phrases.
@@ 5.1.2 The 10,000 dataset
The process of assessing the quality of the hand-coded data was an iterative one. Information about the coding is presented in Table 3. From the final sample, 32 advertisements with a description in a language other than Polish were removed, although they had been coded by experts.
@@
It was observed by experts that for 38 advertisements (0.38%; 13 from the first expert, 13 from the second, and 2 from the third), it was not possible to ascertain the occupation due to a lack of pertinent information or because no occupation in the official classification matched the description. In one instance, the expert indicated that the advertisement describes two distinct occupations. In 59 cases (representing 0.6% of those with codes), experts noted that multiple codes align with the description. Rather than removing these 59 cases, we elected to flag them as potential codes for the same description. Consequently, the final dataset comprised 9,932 ads (prior to accounting for multiple codes).
The subsequent stage of the study involved an investigation into the quality of the coding produced by experts based on a set of 100 ads that exhibited a certain degree of overlap. Two of the experts did not provide a code for one of the advertisements, whereas the others did. Accordingly, for the purposes of comparison, the total number of ads considered was 98. Table 4 presents the data on the rate of agreement of expert coding at the 1-, 2-, 4-, and 6-digit levels (section "before clerical review") along with 95% confidence interval using sampling weights using the survey (Lumley,, 2004) package. In addition, we have estimated Cohen’s kappa coefficient for agreement between 1 digit codes for each pair of experts (we used survey::svykappa() function).
@@
Next, we focus on the assessment of the quality of CBOP coding, as in our sample, we had 236 ads from this source. We were able to link 213 unique ads to the CBOP using the hash identifier present in the ad’s URL. Table 5 presents information on the quality of CBOP coding according to the experts (point and 95% CI for the rate of agreement) and Kappa coefficient for each expert and CBOP as well as all experts (treated as one group) and CBOP as another (the last row denoted as All). Note that due to sparse data, all measures were calculated without the 1 and 6 occupation group for both CBOP and experts (3 obs were dropped).
Table 5: Point and 95%interval estimates of the rate of agreement of expert coding at 1, 2, 4 and 6 digits for CBOP ads
@@
Subsequently, an assessment of discrepancies in coding was conducted by experts. Two experts consented to participate in the study and provided comments. The second expert determined that six of the 29 advertisements provided by CBOP were accurate, differing from the expert’s initial assessment. Additionally, four of the 29 advertisements exhibited a dual accuracy, with both codes being correct. Consequently, the expert’s agreement rate at the 6-digit level would increase to 76%, while at the 1-digit level, it would increase slightly, given that the disagreements are at more detailed levels. The third expert determined that in seven of the 30 advertisements, both codes are accurate. However, the expert highlighted that the lack of sufficient information precludes the ability to distinguish between the two. Consequently, for this expert, the agreement rate at the six-digit level would increase to 72%, while at the one-digit level, a slight increase is observed due to the disagreements occurring at more detailed levels. This also results in lower Kappa coefficients around 75% in comparison to about 85% presented in Table 4 but they are not significantly different if we compare confidence intervals.
Table 6 presents a comparison of agreements at the one-digit level. The figures in the table refer to the precision of the classification of CBOP ads. With the exception of the 4th group, the agreement rate for groups 2 to 9 is above 70%. However, as there is only one example in the first group, we refrain from commenting on it. Overall, it can be observed that the CBOP codes are characterised by a high level of confidence, with the main discrepancies occurring between the second and fourth groups.
Table 6: Comparison of codes at the 1-digit groups between CBOP and experts coding
@@ 5.2 General description of the data
In this section, we present information on the quality of input data. The combined dataset contained 200,875 cases out of which 9,200 cases refer to the official dictionary2020endnote: 20Note that this number is larger than the number of codes because for certain codes more than one case is provided.. Table 7 presents the coverage of occupation 6-digits occupation codes by main groups and source. The last row presents information about the number of rows for a given dataset. Datasets varied in terms of coverage. The most complete was CBOP with coverage close to 80% of all 6-digit codes within given occupation group. Due to sample size, hand-coded data covered only a small fraction of available codes. The combined dataset was missing 186 occupation codes2121endnote: 21Detailed information: 1 group: 9, 2 group: 84, 3 group: 46, 4 group: 1, 5 group: 5, 6 group: 1, 7 group: 27, 8 group: 8, 9 group: 5). List of these codes is provided here https://github.com/OJALAB/job-ads-datasets/blob/main/data/codes-not-coveted.csv.. It should be noted that for the training dataset the complete dictionary with all 2,911 codes was used.
@@
Group Info+ GUS ESCO KPRM Hand (1) Hand (2) CBOP All Cases
@@
Source: Own elaboration. Group (1 digit) as in Table 1.
@@
Source: Own elaboration. Groups as in Table 1.
@@
Sources varied in terms of a number of words describing a given code. Table 8 presents information on the distribution of the number of words in the combined dataset for each group separately. The median number of words was 39 and varied between from 31 for the 9th to 83 for the 1st group. A comparison of the number of words in the descriptions from the two sources reveals that the CBOP descriptions contain, on average, approximately 50 words, while the descriptions from online sources contain more than 120 words on average. The detailed information for each input dataset is presented in the Supplementary Material in Table LABEL:tab-words-by-source.
@@ 1. Step 1:
    Combine datasets: INFO+, KPRM, HAND (10k), HAND (1k), ESCO and CBOP into one dataset.
@@ 4. Step 4:
    Strata with $>1$ cases were sampled using a stratified design with sample sizes defined as in Step 2; 70% for training and 30% for testing.
@@ 5. Step 5:
    The final training dataset was created by combining 70% of the dataset from Step 4 with an fficial dictionary and GUS thesaurus.
@@ 6. Step 6:
Table 9 presents information on the number of cases and 6-digit codes in train and test datasets by source. The training dataset contains close to 143k cases while testing 58k. The main contribution is from the CBOP dataset. The number of codes in the training dataset is equal to the number of codes in the official classification, while in the test data not all codes are available.
@@
In order to verify the approach we translated train and test data from English (based on Google Translate) to 22 languages (all UE languages) with Argos Translate. This dataset over 3,4 mln records for training and close to 1,4 mln for test data. In order to avoid overfitting of the multilingual classifier we selected a stratified random sample of both datasets independently with 10% and 15% from each stratum created by 6-digit occupation code and source type. Information about the number of cases for each dataset (train and test are presented in Table 9 while detailed information about the distribution of languages are presented in Table 10.
@@
Table 10 contains information about the number of tokens used to train XLM-RoBERTa. The primary objective is to demonstrate that some languages are markedly under-represented in the XLM-RoBERTa dataset, including Estonian, Irish, Swedish and Ukrainian, while others are over-represented, such as German and Russian. This may suggest that the quality of classification may vary between languages.
@@ 6.1 Models and implementation
In the empirical study, we considered the following model variants of bottom-up and top-down approaches described in Section 3.2:
@@
For the implementation of the Transformer architecture, we use PyTorch (Paszke et al.,, 2019), PyTorch Lightning (Falcon and Team,, 2019), and Huggingface Transformers (Wolf,, 2019) libraries. We fine-tune transformer-based models using AdamW optimizer  (Loshchilov and Hutter,, 2017), with the base learning rate equal to $51\mathrm{e-}$ 5, weight decay to $0.01$ (for all layers except bias terms and layer norms layers), and Adam hyperparameters to $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=1\mathrm{e-}8$. We use a linear learning rate schedule with a warm-up phase of 50 training steps. The training was done on a machine with two Nvidia RTX5000 GPUs with 24 GB RAM each. For base models, we used 10 epochs, while for large, only 5. The reason for that was the computational time, for instance, for HerBERT-base one epoch was about 30 min, for XLM-RoBERTa-base about an hour, for HerBERT-large about 1.5 hours and XLM-RoBERTa-large about 2.5 hours. Training for the top-down approach was approximately two times longer than for the bottom-up approach.
@@ 6.2 Evaluation metrics
To evaluate the quality of obtained classifiers, we primarily report a per hierarchy level logistic loss (cross-entropy) to investigate if the proposed approach improves the quality of obtained estimates of conditional probabilities of classes $\eta_{c}(\boldsymbol{x}):=\mathbb{P}(y_{c}=1|\boldsymbol{x})$ at each level of the hierarchy. Additionally, we report recall@$k$ at each level of the hierarchy. which indicates how many of all relevant classes were predicted in the set of $k$ predictions coming from a classifier. Recall is formally defined as:
$\operatorname{recall}@k(\boldsymbol{y},\boldsymbol{h}(x))=\frac{1}{||\boldsymbol{y}||_{1}}\sum_{j\in\hat{\mathcal{Y}}_{k}}\llbracket y_{j}=1\rrbracket$ (9)
where $||\boldsymbol{y}||_{1}$ is $L_{1}$ norm of vector $\boldsymbol{y}$ – number of positive classes, $\hat{\mathcal{Y}}_{k}$ is a set of $k$ labels with top conditional probabilities $\eta_{c}(\boldsymbol{x})$ predicted by $\boldsymbol{h}$ for $\boldsymbol{x}$. We calculate recall@$k$ separately for each level. Notice that in this case, $||\boldsymbol{y}||_{1}$ is always equal to 1, as only one class is relevant per hierarchy level. Because of that, in this case, recall@$1$ is indifferent to accuracy. It is also easy to see that the optimal decision rule when predicting for recall@$k$ is to select top-$k$ classes with the highest conditional probabilities $\eta_{c}(\boldsymbol{x})$ (Lapin et al.,, 2018). This makes recall@$k$ align with our goal of measuring the quality of probability estimates while being more intuitive than logistic loss.

### v2

@@ 4.3.1 The 10,000 ads dataset
Before sampling, the data was cleaned to remove rows that had no job description (233,464 observations), had zero words or less than 5 characters1818endnote: 18The number of words was counted using stri_count_words function from the stringi package and number of characters were counted using nchar from the base R (93 observations). Then we removed records with non-Polish descriptions. We have used Google’s Compact Language Detector 3 available through the cld3 package in R (Ooms,, 2024) to detect the language of the job description and identified 181,425 rows with non-Polish text. In the end, the final dataset used for sampling consisted of 1,390,985 unlabelled records (77% of the initial dataset).
@@
The second stratification variable was based on the number of words in the description that were categorized into the following groups: $[1,10]$, $(10,25]$, $(25,50]$, $(50,100]$, $(100,150]$, $(150,200]$, $(200,300]$ and $300+$. Table˜14 presents information about stratification and population sizes within these strata. The largest strata was "specjaliści" (specialists) with over 130k ads. The second largest were "sprzedawcy" (sellers) with close to 55k ads. The smallest group were "sprzątacze" (cleaners) with over 7k ads. According to the second stratification variable groups with over 50 words were more or less of the same size. There were about 40k ads with at most 10 words and 70k with at most 25 words.
In order to code the dataset, three experts from PEOs were recruited in February 2022. Due to budgetary limitations, the sample size was set at 10,000, with each expert responsible for coding 3,333 advertisements. However, due to rounding, the final sample size was 10,002. The sample size was selected using a stratified systematic sampling design, with probabilities proportional to the number of characters in a given ad. The sample was selected using strata from the sampling package (Tillé and Matei,, 2023). Further details regarding the sample size according to the source can be found in Table˜15. The primary source was pracuj.pl, praca.egospodarka and aplikuj.pl. It should be noted that 241 ads were from ePraca, which enabled an assessment of the quality of ePraca encoding.
A total of 3,233 advertisements were selected for each expert, with an additional 100 advertisements sampled for cross-validation of expert coding1919endnote: 19it should be noted that for 4 ads we also had information from ePraca. Each expert was given an Excel spreadsheet, available via Google Sheets, containing all the relevant information. The experts were tasked with coding the advertisements into six-digit codes and providing comments where necessary. They were given a one-month period to complete the coding of the advertisements into occupation codes.
@@ 4.3.2 The 1,000 ads dataset
In January 2023, an additional dataset of 1,000 advertisements was selected for manual coding by a single expert. The sample was selected from a comprehensive dataset of over 543,000 ads scraped from one of the largest services in Poland. In the case of this particular dataset, we were able to identify in excess of 227,000 unique descriptions. We then proceeded to implement a sample selection process in accordance with the following scheme: the specification of a search phrase (including regular expression notation) and the subsequent selection of the most appropriate samples. The information pertaining to the aforementioned phrases, the selected advertisements and the number of cases in one of the largest services in Poland is detailed in Table˜16.
@@ 4.4 The multilingual dataset
Firstly, the Polish text was translated into English using Google Translate. This was achieved by utilising the Google Sheets application and the googletranslate() function. This approach was employed for the translation of over 200,000 examples. Additionally, Argos Translate was employed for the translation of the remaining text. The open-source offline translation library, written in Python, is available at https://github.com/argosopentech/argos-translate. It uses the OpenNMT library (Klein et al.,, 2017), which in turn uses the CTranslate2 library, which is a fast inference engine for Transformer models. The objective was to evaluate the influence of an automated language library on the classification of a bilingual dataset2020endnote: 20Unfortunately, information on the quality of the automatic translation is rather limited. For instance, Klein et al., (2020) reports that the OpenNMT/CTranslate2 library scored 26.7% in the Bilingual Evaluation Understudy (BLEU) task (see Papineni et al., (2002)) for English to German translation. Author of the Argos Translate fine-tuned the OpenNMT model in the OPUS – A Collection of Multilingual Parallel Corpora with Tools and Interfaces (see Tiedemann, (2009)). Certainly, the quality of automated translation has an effect on the classifier, as we show in the comparison between the Polish-English translations using Google Translate and Argos Translate..
@@ 5.1.1 The 1,000 dataset
The 1,000 dataset was coded by a single expert, and thus, cross-validation was not available. The expert identified 968 advertisements for which a single occupation code was deemed appropriate, while 32 advertisements were assigned multiple codes (27 with two codes, three with three codes, and one with four codes). Therefore, the final dataset comprised 1,035 advertisements with 226 codes, with the most prevalent being: "Programista aplikacji" (Applications programmer, 125 occurrences; 251401), "Analityk baz danych (data scientist)" (Data scientist, 77 occurrences; 252102) and "Specjalista do spraw cyberbezpieczeństwa" (Cybersecurity specialist, 63 occurrences; 252902). See Table˜17 for the information on the number of unique 6-digit and 1-digit codes for a given stratum defined by the search phrases.
@@ 5.1.2 The 10,000 dataset
The process of assessing the quality of the hand-coded data was an iterative one. Information about the coding is presented in Table˜3. From the final sample, 32 advertisements with a description in a language other than Polish were removed, although they had been coded by experts.
@@
It was observed by experts that for 38 advertisements (0.38%; 13 from the first expert, 13 from the second, and 2 from the third), it was not possible to ascertain the occupation due to a lack of pertinent information or because no occupation in the official classification matched the description. In one instance, the expert indicated that the advertisement describes two distinct occupations. In 59 cases (representing 0.6% of those with codes), experts noted that multiple codes align with the description. Rather than removing these 59 cases, we selected to flag them as potential codes for the same description. Consequently, the final dataset comprised 9,932 ads (prior to accounting for multiple codes).
The subsequent stage of the study involved an investigation into the quality of the coding produced by experts based on a set of 100 ads that exhibited a certain degree of overlap. Two of the experts did not provide a code for one of the advertisements, whereas the others did. Accordingly, for the purposes of comparison, the total number of ads considered was 98. Table˜4 presents the data on the rate of agreement of expert coding at the 1-, 2-, 4-, and 6-digit levels (section "before clerical review") along with 95% confidence interval using sampling weights using the survey (Lumley,, 2004) package. In addition, we have estimated Cohen’s kappa coefficient for agreement between 1 digit codes for each pair of experts (we used survey::svykappa() function).
@@
Next, we focus on the assessment of the quality of ePraca coding, as in our sample, we had 236 ads from this source. We were able to link 213 unique ads to the ePraca using the hash identifier present in the ad’s URL. Table˜5 presents information on the quality of ePraca coding according to the experts (point and 95% CI for the rate of agreement) and Kappa coefficient for each expert and ePraca as well as all experts (treated as one group) and ePraca as another (the last row denoted as All). Note that due to sparse data, all measures were calculated without the 1 and 6 occupation group for both ePraca and experts (3 obs were dropped).
Table 5: Point and 95% interval estimates of the rate of agreement of expert coding at 1, 2, 4 and 6 digits for ePraca ads
@@
Subsequently, an assessment of discrepancies in coding was conducted by experts. Two experts consented to participate in the study and provided comments. The second expert determined that six of the 29 advertisements provided by ePraca were accurate, differing from the expert’s initial assessment. Additionally, four of the 29 advertisements exhibited a dual accuracy, with both codes being correct. Consequently, the expert’s agreement rate at the 6-digit level would increase to 76%, while at the 1-digit level, it would increase slightly, given that the disagreements are at more detailed levels. The third expert determined that in seven of the 30 advertisements, both codes are accurate. However, the expert highlighted that the lack of sufficient information precludes the ability to distinguish between the two. Consequently, for this expert, the agreement rate at the six-digit level would increase to 72%, while at the one-digit level, a slight increase is observed due to the disagreements occurring at more detailed levels. This also results in lower Kappa coefficients around 75% in comparison to about 85% presented in Table˜4 but they are not significantly different if we compare confidence intervals.
Table˜6 presents a comparison of agreements at the one-digit level. The figures in the table refer to the precision of the classification of ePraca ads. With the exception of the 4th group, the agreement rate for groups 2 to 9 is above 70%. However, as there is only one example in the first group, we refrain from commenting on it. Overall, it can be observed that the ePraca codes are characterised by a high level of confidence, with the main discrepancies occurring between the second and fourth groups.
Table 6: Comparison of codes at the 1-digit groups between ePraca and experts coding
@@ 5.2 General description of the data
In this section, we present information on the quality of input data. The combined dataset contained 200,875 cases out of which 9,200 cases refer to the official dictionary2121endnote: 21Note that this number is larger than the number of codes because for certain codes more than one case is provided.. Table˜7 presents the coverage of occupation 6-digits occupation codes by main groups and source. The last row presents information about the number of rows for a given dataset. Datasets varied in terms of coverage. The most complete was ePraca with coverage close to 80% of all 6-digit codes within given occupation group. Due to sample size, hand-coded data covered only a small fraction of available codes. The combined dataset was missing 186 occupation codes2222endnote: 22Detailed information: 1 group: 9, 2 group: 84, 3 group: 46, 4 group: 1, 5 group: 5, 6 group: 1, 7 group: 27, 8 group: 8, 9 group: 5). List of these codes is provided here https://github.com/OJALAB/job-ads-datasets/blob/main/data/codes-not-coveted.csv.. It should be noted that for the training dataset the complete dictionary with all 2,911 codes was used.
@@
Group Info+ GUS ESCO KPRM Hand (1) Hand (2) ePraca All Cases
@@
Source: Own elaboration. Group (1 digit) as in Table˜1.
@@
Source: Own elaboration. Groups as in Table˜1.
@@
Sources varied in terms of a number of words describing a given code. Table˜8 presents information on the distribution of the number of words in the combined dataset for each group separately. The median number of words was 39 and varied between from 31 for the 9th to 83 for the 1st group. A comparison of the number of words in the descriptions from the two sources reveals that the ePraca descriptions contain, on average, approximately 50 words, while the descriptions from online sources contain more than 120 words on average. The detailed information for each input dataset is presented in the Supplementary Material in Table LABEL:tab-words-by-source.
@@ 1. Step 1:
    Combine datasets: INFO+, KPRM, HAND (10k), HAND (1k), ESCO and ePraca into one dataset.
@@ 4. Step 4:
    Strata with >1>1 cases were sampled using a stratified design with sample sizes defined as in Step 2; 70% for training and 30% for testing.
@@ 5. Step 5:
    The final training dataset was created by combining 70% of the dataset from Step 4 with an official dictionary and GUS thesaurus.
@@ 6. Step 6:
Table˜9 presents information on the number of cases and 6-digit codes in train and test datasets by source. The training dataset contains close to 143k cases while testing 58k. The main contribution is from the ePraca dataset. The number of codes in the training dataset is equal to the number of codes in the official classification, while in the test data not all codes are available.
@@
In order to verify the approach we translated train and test data from English (based on Google Translate) to 22 languages (all UE languages) with Argos Translate. This dataset over 3,4 mln records for training and close to 1,4 mln for test data. In order to avoid overfitting of the multilingual classifier we selected a stratified random sample of both datasets independently with 10% and 15% from each stratum created by 6-digit occupation code and source type. Information about the number of cases for each dataset (train and test are presented in Table˜9 while detailed information about the distribution of languages are presented in Table˜10.
@@
Table˜10 contains information about the number of tokens used to train XLM-RoBERTa. The primary objective is to demonstrate that some languages are markedly under-represented in the XLM-RoBERTa dataset, including Estonian, Irish, Swedish and Ukrainian, while others are over-represented, such as German and Russian. This may suggest that the quality of classification may vary between languages.
@@ 6.1 Models and implementation
In the empirical study, we considered the following model variants of bottom-up and top-down approaches described in Section˜3.2:
@@
For the implementation of the Transformer architecture, we use PyTorch (Paszke et al.,, 2019), PyTorch Lightning (Falcon and Team,, 2019), and Huggingface Transformers (Wolf et al.,, 2020) libraries. We fine-tune transformer-based models using AdamW optimizer  (Loshchilov and Hutter,, 2019), with the base learning rate equal to $1\mathrm{e-}5$, weight decay to $0.01$ (for all layers except bias terms and layer norms layers), and Adam hyperparameters to $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=1\mathrm{e-}8$. We use a linear learning rate schedule with a warm-up phase of 50 training steps. The training was done on a machine with two Nvidia RTX5000 GPUs with 24 GB RAM each. For base models, we used 10 epochs, while for large, only 5. The reason for that was the computational time, for instance, for HerBERT-base one epoch was about 30 min, for XLM-RoBERTa-base about an hour, for HerBERT-large about 1.5 hours and XLM-RoBERTa-large about 2.5 hours. Training for the top-down approach was approximately two times longer than for the bottom-up approach.
@@ 6.2 Evaluation metrics
To evaluate the quality of obtained classifiers, we primarily report a per hierarchy level logistic loss (cross-entropy) to investigate if the proposed approach improves the quality of obtained estimates of conditional probabilities of classes $\eta_{c}(\boldsymbol{x})\coloneqq\mathbb{P}(y_{c}=1|\boldsymbol{x})$ at each level of the hierarchy. Additionally, we report recall@k@k at each level of the hierarchy. which indicates how many of all relevant classes were predicted in the set of $k$ predictions coming from a classifier. Recall is formally defined as:
[…]
where $\boldsymbol{y}||_{1}$ is $L_{1}$ norm of vector $\boldsymbol{y}$ – number of positive classes, $\hat{\mathcal{Y}}_{k}$ is a set of $k$ labels with top conditional probabilities $\eta_{c}(\boldsymbol{x})$ predicted by $\boldsymbol{h}$ for $\boldsymbol{x}$. We calculate recall@k@k separately for each level. Notice that in this case, $\boldsymbol{y}||_{1}$ is always equal to 1, as only one class is relevant per hierarchy level. Because of that, in this case, recall@1@1 is indifferent to accuracy. It is also easy to see that the optimal decision rule when predicting for recall@k@k is to select top-$k$ classes with the highest conditional probabilities $\eta_{c}(\boldsymbol{x})$ (Lapin et al.,, 2018). This makes recall@k@k align with our goal of measuring the quality of probability estimates while being more intuitive than logistic loss.